# The Low Energy Limit of BFSS Quantum Mechanics

### v1

@@ Introduction.
Shortly after [1], Ref. [5] proposed a network of additional holographic dualities pairing maximally supersymmetric theories with their gravitational duals. These were identified by studying D-branes in string theory and decoupling the brane worldvolume theory from gravity. This letter focuses on the simplest duality of [5], namely, between the low-energy limit of D00-branes and maximally supersymmetric quantum mechanics (BFSS) [6]. Recent years have seen the development of novel techniques for analyzing the gauge theory side of this duality at strong coupling using computer lattice simulations [7, 8, 9, 10, 11, 12, 13, 14, 15]. Our study will offer key insights into the gravitational aspect of the duality, crucial for understanding BFSS theory in its strongly coupled regime and enabling precise holography matchings.
@@
For sufficiently high temperatures, namely when $N^{-5/6}\lesssim\tau\ll 1$, [5] conjectured that the dominant phase retains translational invariance in $z$, has horizon topology $S^{1}\times S^{8}$, and is simply given by
@@
However, for sufficiently low temperatures, specifically $N^{-25/21}\ll\tau\lesssim N^{-5/6}$, [5] conjectured that the relevant phase should be a localized black hole on the M-theory circle. In this letter, we will show that while this expectation is met in the canonical ensemble, it surprisingly fails in the microcanonical ensemble. Indeed, we will present compelling numerical evidence for the existence of a dominant phase, within a specific range of energies, that breaks translational invariance in $z$ while maintaining a horizon topology of $S^{1}\times S^{8}$. Morover, in the canonical ensemble, we will determine the precise temperature at which the first-order transition occurs between localised black holes and the translationally preserving phase.
@@
where $\Gamma^{a}_{cd}(\mathfrak{g})$ is the Christoffel connection associated to a metric $\mathfrak{g}$, and $\bar{g}$ is a reference metric which we are free to choose. Solutions to the Einstein-DeTurck equation are generally not always Ricci-flat. However, if we focus on static configurations that admit a regular Euclidean section, and have Kaluza-Klein asymptotics, it has been shown in [19] that solutions with $\xi\neq 0$ do not exist. As such, for such spacetimes, we expect solutions of the Einstein-DeTurck equation to coincide with Ricci-flat solutions. Solving (7) instead of $R_{ab}=0$ offers significant advantages. Once a reference metric $\bar{g}$ is chosen, the Einstein-DeTurck equation forms a well-posed elliptic system [19]. Solutions must have $\xi^{a}=0$, implying $\Box x^{a}=g^{cd}\Gamma^{a}_{cd}(\bar{g})$. Thus, these solutions are naturally expressed in generalized harmonic coordinates [20], with the gauge source term fully controlled by $\bar{g}$. A good gauge choice involves choosing $\bar{g}$ judiciously. The choices of $\bar{g}$ across the solutions described in this letter are presented in the Supplemental Material.
@@
$+L^{2}Q_{4}(X,Z)\left[\mathrm{d}Z+Q_{3}(X,Z){\rm d}X\right]^{2}+\frac{Q_{5}(X,Z)}{(1-X^{2})^{2}}\mathrm{d}\Omega_{8}^{2}\,,$
@@
As an example, the Carrollian transformation (13), takes the uniform black string (11) to (6), which we recall describes a finite temperature stack of D0 branes. Furthermore, since the static Kaluza-Klein circle has size $L$ at spatial infinity, and we require the M-theory circle -- parametrised by $z$- to have size $L_{11}=2\pi g_{s}\ell_{s}$ there, the transformation (13) uniquely determines $L$ to be:
@@ Results.
We begin with the canonical ensemble, fixing the temperature $\tau N^{5/9}$ and identifying the phase with the lower free energy density, $\mathfrak{f}N^{-4/9}$. Fig. 2 shows $\Delta\mathfrak{f}$, the free energy difference between a given BFSS thermal phase and the uniform phase at the same $\tau$. The black disks show exact numerical data for the localized BFSS phase [^2] In [42], we compare our numerical localized solutions with the second-order perturbative results from [43] and find excellent agreement in regions of moduli space where localized black holes do exist., and the red squares correspond to the non-uniform BFSS phase. We observe a first order phase transition between the localised phase and the uniform phase at $\tau N^{5/9}\approx 0.6118$. This is one of the main results of this letter. This aligns with the phase diagram conjectured by [5], but here we provide a precise prediction for the critical temperature where the first-order phase transition occurs. In principle, it should be possible to derive the latter from a field theory calculation.
@@
On the other hand, the microcanonical ensemble reveals a much more interesting and unexpected structure. In Fig. 3, using the same color coding as Fig. 2, we plot $\Delta\sigma/N$, the entropy density difference between a given BFSS phase and the uniform phase at fixed energy density $\varepsilon N^{-4/9}$. For $0<\varepsilon<\varepsilon_{\rm m}$, the localized (black) phase dominates, while for $\varepsilon_{\rm m}<\varepsilon<\varepsilon_{\rm M}$, the non-uniform (red) phase dominates the microcanonical ensemble. A topology-changing transition occurs at $\varepsilon=\varepsilon_{\rm m}$, as described by Kol in [44]. One finds that $1.505(4)\lesssim\varepsilon_{\rm m}\lesssim 1.512(6)$, where the lower bound is found computing the maximum value of $\varepsilon$ for which localized solutions exist, and the upper bound follows from finding the minimum value of $\varepsilon$ for which non-uniform solutions exist. Improved numerics would pin down $\varepsilon_{\rm m}$ even further. On the other hand, we find the value of $\varepsilon_{\rm M}$ to be $\varepsilon_{\rm M}\approx 1.713(2)$. This simply follows from applying the map (15) to the gravitational Gregory-Laflamme instability onset of the uniform phase. It remains a formidable challenge to predict $\varepsilon_{\rm m}$ and $\varepsilon_{\rm M}$ from a field theory calculation in the microcanonical ensemble and we believe our precise identification of these values should motivate and guide such a search. The existence of an energy window where the non-uniform phase dominates the microcanonical ensemble is surprising and constitutes the main result of this letter.
@@ Discussion.
While there has been extensive research in high-energy physics aimed at understanding the canonical ensemble of various field theories -- using approaches such as localization, see for instance [47] and references therein -- it is only recently that the microcanonical ensemble has garnered significant attention [48, 49, 50]. The increased complexity of studying the microcanonical ensemble stems from the necessity of understanding the detailed nature of field theory microstates. Addressing this challenge is notoriously difficult and is a key reason why a complete solution to the black hole information paradox remains elusive. Indeed, recent calculations show how to compute the entropy of Hawking radiation (see for instance [51, 52, 53, 54, 55]), but not how to compute its precise quantum state [56]. Our work reveals an intricate microcanonical ensemble structure for BFSS quantum mechanics at low energies. It would be desirable to identify all the relevant phases using BFSS degrees of freedom directly, though this task will invariantly require an understanding of the underlying microstates.
@@
The Carrollian map (13) together with the associated thermodynamical map (15) allows us to associate any black hole solution with standard Kaluza-Klein asymptotics into one with D0 brane asymptotics (and thus with any BFSS state). In particular, given the multitude of black hole solutions in higher dimensions -- see, for instance, [57, 58, 59, 60, 61] -- and the nontrivial nature of the map (15), one wonders how many more unexpected phases can possibly dominate a generalized ensemble that includes rotations and/or supersymmetric mass deformations à la BMN [62, 63]. We leave these exploratory investigations to future work.
@@ Appendix A Constructing the several black hole phases
We use the Einstein-de Turck formulation of the gravitational equations of motion [16, 19, 18, 17]. This formulation requires that we first choose a reference metric $\overline{g}$. This metric need not be a solution to the Einstein equation, but must contain the same symmetries and causal structure as the desired solution and, to facilitate the numerical search, should resemble the final solution near the horizon and asymptotically. With the reference metric chosen, the DeTurck method then modifies the Einstein equation $R_{\mu\nu}=0$ to (7)-(LABEL:eq:ede2). Unlike $R_{\mu\nu}=0$, the Einstein-de Turck equation yields PDEs that are elliptic in character and solutions of the Einstein-de Turck system are also solutions of 11-dimensional vacuum Einstein gravity as long as the de Turck vector vanishes, $\xi_{\mu}\xi^{\mu}=0$, which is what we want.
@@ A.1 Localized black holes
Figure 5: Integration domain with two patches $I$ and $II$. Chebyshev-Gauss-Lobatto grids with $50\times 50$ points are placed using transfinite interpolation. Left panel: Patch $I$ (in the far region) uses $\{\rho,\xi\}$ coordinates and patch $II$ (near the horizon in the quarter circle in the lower left) is mapped from $\{x,y\}$ coordinates using (A.1). Right panel: Patch II (near the horizon) uses $\{x,y\}$ coordinates.
The Einstein-de Turck equations motions -- (7) in the main text -- for (A.1) must be solved together with the boundary conditions that we now discuss. At the asymptotic boundary $\rho=1$ we impose Dirichlet conditions ${\tilde{f}}_{1,2,3,4}=1,{\tilde{f}}_{5}=0$: the solution must approach the reference metric. At the horizon $y=1$, regularity of the solution requires that $f_{i}$ obey certain Robin boundary conditions (whose expressions are long and unilluminating). At the remaining boundaries, $f_{i}$ and $\tilde{f}_{i}$ must obey Neumann conditions either due to regularity at the axis $\rho=0$ (or $x=1$), or reflection symmetry at $\xi=0$ (or $x=0$) and at $\xi=1$ (or $y=0$).
@@
Localized black holes are a one-parameter family of solutions that we can take to be parametrized by the temperature $\mathcal{T}^{(0)}_{H}$, that is to say, by $y_{0}$ for our numerical solutions. $L$ just sets a scale and it drops out of the equations of motion. As a first seed for the Newton-Raphson algorithm, we choose the reference metric with $y_{0}=1/10$ and then march our code for increasing values of $y_{0}$. Our selection of reference metric has restricted our temperature range to $T\,L>3/2^{4/3}\approx 1.19$; this is (A.10) with $y_{0}=0$. We are actually not far from saturating this bound (our minimum $y_{0}$ is 0.1). Should we wish to pin down more accurately where the localized black hole terminates and merges through a topology-changing transition [44] to the non-uniform black string we could first decrease $y_{0}$ further to approach 0. Then, if necessary we could still change our reference metric to permit even lower temperatures. To get the Helmoltz free energy $\mathcal{F}^{(0)}$ we integrate the first law of thermodynamics, $\mathrm{d}\mathcal{F}^{(0)}=-\mathcal{S}^{(0)}_{H}\,\mathrm{d}\mathcal{T}^{(0)}_{H}$, and the energy is then $\mathcal{E}^{(0)}=\mathcal{F}^{(0)}+\mathcal{T}^{(0)}_{H}\mathcal{S}^{(0)}_{H}$. The tension $\mathcal{T}_{Z}^{(0)}$ along the circle $S^{1}_{L}$ can be obtained from the Smarr relation (12), $8\,\mathcal{E}^{(0)}=9\,\mathcal{T}_{H}^{(0)}\,\mathcal{S}_{H}^{(0)}+\mathcal{T}_{Z}^{(0)}$. The values of these thermodynamics quantities are in excellent agreement with the results obtained using the standadard Arnowitt-Deser-Misner formalism [43, 65] or the covariant Noether charge formalism (a.k.a. covariant phase space method) [40, 41].
@@ A.2 Non-uniform strings
where the $\ldots$ has two types of contributions. First, we have higher order polynomial contributions of the form $(1-x)^{n}$ with $n\geq 10$. Second, there are nonperturbative terms in $1-x$ of the form $e^{-\frac{2\pi\,n}{1-x}}\cos(2\pi ny)$ (for $i=1,2,4,5$) and $e^{-\frac{2\pi\,n}{1-x}}\sin(2\pi ny)$ (for $i=3$). The coefficients $\delta_{i}$ are independent of $y$ and will later parametrise the energy and tension of the non-uniform string.

### v2

@@ Introduction.
Shortly after [1], Ref. [7] proposed a network of additional holographic dualities pairing maximally supersymmetric theories with their gravitational duals (see also [8, 9, 10]). These were identified by studying D-branes in string theory and decoupling the brane worldvolume theory from gravity. This letter focuses on the simplest duality of [7], namely, between the low-energy limit of D0-branes and maximally supersymmetric quantum mechanics (BFSS) [11]. Recent years have seen the development of novel techniques for analyzing the gauge theory side of this duality at strong coupling using computer lattice simulations [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 5, 22]. Our study will offer key insights into the gravitational aspect of the duality, crucial for understanding BFSS theory in its strongly coupled regime and enabling precise holography matchings.
@@
For sufficiently high temperatures, namely when $N^{-5/9}\lesssim\tau\ll 1$, [7] conjectured that the dominant phase retains translational invariance in $z$, has horizon topology $S^{1}\times S^{8}$, and is simply given by
@@
However, for sufficiently low temperatures, specifically $N^{-5/6}\ll\tau\lesssim N^{-5/9}$, [7] conjectured that the relevant phase should be a localized black hole on the M-theory circle. In this letter, we will show that while this expectation is met in the canonical ensemble, it surprisingly fails in the microcanonical ensemble. Indeed, we will present compelling numerical evidence for the existence of a dominant phase, within a specific range of energies, that breaks translational invariance in $z$ while maintaining a horizon topology of $S^{1}\times S^{8}$. Morover, in the canonical ensemble, we will determine the precise temperature at which the first-order transition occurs between localised black holes and the translationally preserving phase.
@@
where $\Gamma^{a}_{cd}(\mathfrak{g})$ is the Christoffel connection associated to a metric $\mathfrak{g}$, and $\bar{g}$ is a reference metric which we are free to choose. Solutions to the Einstein-DeTurck equation are generally not always Ricci-flat. However, if we focus on static configurations that admit a regular Euclidean section, and have Kaluza-Klein asymptotics, it has been shown in [26] that solutions with $\xi\neq 0$ do not exist. As such, for such spacetimes, we expect solutions of the Einstein-DeTurck equation to coincide with Ricci-flat solutions. Solving (7) instead of $R_{ab}=0$ offers significant advantages. Once a reference metric $\bar{g}$ is chosen, the Einstein-DeTurck equation forms a well-posed elliptic system [26]. Solutions must have $\xi^{a}=0$, implying $\Box x^{a}=g^{cd}\Gamma^{a}_{cd}(\bar{g})$. Thus, these solutions are naturally expressed in generalized harmonic coordinates [cmp/1104114004], with the gauge source term fully controlled by $\bar{g}$. A good gauge choice involves choosing $\bar{g}$ judiciously. The choices of $\bar{g}$ across the solutions described in this letter are presented in the Supplemental Material.
@@
$\hskip-11.38092pt+L^{2}Q_{4}(X,Z)\left[\mathrm{d}Z+Q_{3}(X,Z){\rm d}X\right]^{2}+\frac{Q_{5}(X,Z)}{(1-X^{2})^{2}}\mathrm{d}\Omega_{8}^{2}\,$,
@@
As an example, the Carrollian transformation (13), takes the uniform black string (11) to (6), which we recall describes a finite temperature stack of D0 branes. Furthermore, since the static Kaluza-Klein circle has size $L$ at spatial infinity, and we require the M-theory circle - parametrised by $z$ - to have size $L_{11}=2\pi g_{s}\ell_{s}$ there, the transformation (13) uniquely determines $L$ to be:
@@ Results.
We begin with the canonical ensemble, fixing the temperature $\tau N^{5/9}$ and identifying the phase with the lower free energy density, $\mathfrak{f}N^{-4/9}$. Fig. 2 shows $\Delta\mathfrak{f}$, the free energy difference between a given BFSS thermal phase and the uniform phase at the same $\tau$. The black disks show exact numerical data for the localized BFSS phase [^2] In a forthcoming publication [48], we compare our numerical localized solutions with the second-order perturbative results from [49] and find excellent agreement in regions of moduli space where localized black holes do exist. , and the red squares correspond to the non-uniform BFSS phase. We observe a first order phase transition between the localised phase and the uniform phase at $\tau N^{5/9}\approx 0.6118$. This is one of the main results of this letter. This aligns with the phase diagram conjectured by [7], but here we provide a precise prediction for the critical temperature where the first-order phase transition occurs. In principle, it should be possible to derive the latter from a field theory calculation. We have also computed the specific heat associated with all the dominant phases in the canonical ensemble and found it to be always positive. At first glance, this might seem at odds with expectations for small localized black holes. However, the map given in (15) ensures that the specific heat remains positive even for the localized black hole phase.
@@
On the other hand, the microcanonical ensemble reveals a much more interesting and unexpected structure. In Fig. 3, using the same color coding as Fig. 2, we plot $\Delta\sigma/N$, the entropy density difference between a given BFSS phase and the uniform phase at fixed energy density $\varepsilon N^{-4/9}$. For $0<\varepsilon<\varepsilon_{\rm m}$, the localized (black) phase dominates, while for $\varepsilon_{\rm m}<\varepsilon<\varepsilon_{\rm M}$, the non-uniform (red) phase dominates the microcanonical ensemble. A topology-changing transition occurs at $\varepsilon=\varepsilon_{\rm m}$, as described by Kol in [50]. One finds that $1.505(4)\lesssim\varepsilon_{\rm m}N^{-4/9}\lesssim 1.512(6)$, where the lower bound is found computing the maximum value of $\varepsilon$ for which localized solutions exist, and the upper bound follows from finding the minimum value of $\varepsilon$ for which non-uniform solutions exist. Improved numerics would pin down $\varepsilon_{\rm m}$ even further. On the other hand, we find the value of $\varepsilon_{\rm M}$ to be $\varepsilon_{\rm M}N^{-4/9}\approx 1.713(2)$. This simply follows from applying the map (15) to the gravitational Gregory-Laflamme instability onset of the uniform phase. It remains a formidable challenge to predict $\varepsilon_{\rm m}$ and $\varepsilon_{\rm M}$ from a field theory calculation in the microcanonical ensemble and we believe our precise identification of these values should motivate and guide such a search. The existence of an energy window where the non-uniform phase dominates the microcanonical ensemble is surprising and constitutes the main result of this letter.
@@ Discussion.
While there has been extensive research in high-energy physics aimed at understanding the canonical ensemble of various field theories - using approaches such as localization, see for instance [53] and references therein - it is only recently that the microcanonical ensemble has garnered significant attention [54, 55, 56]. The increased complexity of studying the microcanonical ensemble stems from the necessity of understanding the detailed nature of field theory microstates. Addressing this challenge is notoriously difficult and is a key reason why a complete solution to the black hole information paradox remains elusive. Indeed, recent calculations show how to compute the entropy of Hawking radiation (see for instance [57, 58, 59, 60, 61]), but not how to compute its precise quantum state [62]. Our work reveals an intricate microcanonical ensemble structure for BFSS quantum mechanics at low energies. It would be desirable to identify all the relevant phases using BFSS degrees of freedom directly, though this task will invariantly require an understanding of the underlying microstates.
@@
The Carrollian map (13) together with the associated thermodynamical map (15) allows us to associate any black hole solution with standard Kaluza-Klein asymptotics into one with D0 brane asymptotics (and thus with any BFSS state). In particular, given the multitude of black hole solutions in higher dimensions - see, for instance, [63, 64, 65, 66, 67] - and the nontrivial nature of the map (15), one wonders how many more unexpected phases can possibly dominate a generalized ensemble that includes rotations and/or supersymmetric mass deformations à la BMN [68, 69]. We leave these exploratory investigations to future work.
@@ Appendix A Constructing the several black hole phases
We use the Einstein-de Turck formulation of the gravitational equations of motion [23, 26, 25, 24]. This formulation requires that we first choose a reference metric $\overline{g}$. This metric need not be a solution to the Einstein equation, but must contain the same symmetries and causal structure as the desired solution and, to facilitate the numerical search, should resemble the final solution near the horizon and asymptotically. With the reference metric chosen, the DeTurck method then modifies the Einstein equation $R_{\mu\nu}=0$ to (7)-(8). Unlike $R_{\mu\nu}=0$, the Einstein-de Turck equation yields PDEs that are elliptic in character and solutions of the Einstein-de Turck system are also solutions of 11-dimensional vacuum Einstein gravity as long as the de Turck vector vanishes, $\xi_{\mu}\xi^{\mu}=0$, which is what we want.
@@ A.1 Localized black holes
Figure 5: Integration domain with two patches $I$ and $II$. Chebyshev-Gauss-Lobatto grids with $50\times 50$ points are placed using transfinite interpolation. Left panel: Patch II (in the far region) uses $\{\rho,\xi\}$ coordinates and patch IIII (near the horizon in the quarter circle in the lower left) is mapped from $\{x,y\}$ coordinates using (A.1). Right panel: Patch II (near the horizon) uses $\{x,y\}$ coordinates.
The Einstein-de Turck equations motions - (7) in the main text - for (A.1) must be solved together with the boundary conditions that we now discuss. At the asymptotic boundary $\rho=1$ we impose Dirichlet conditions ${\tilde{f}}_{1,2,3,4}=1,{\tilde{f}}_{5}=0$: the solution must approach the reference metric. At the horizon $y=1$, regularity of the solution requires that $f_{i}$ obey certain Robin boundary conditions (whose expressions are long and unilluminating). At the remaining boundaries, $f_{i}$ and $\tilde{f}_{i}$ must obey Neumann conditions either due to regularity at the axis $\rho=0$ (or $x=1$), or reflection symmetry at $\xi=0$ (or $x=0$) and at $\xi=1$ (or $y=0$).
@@
Localized black holes are a one-parameter family of solutions that we can take to be parametrized by the temperature $\mathcal{T}^{(0)}_{H}$, that is to say, by $y_{0}$ for our numerical solutions. $L$ just sets a scale and it drops out of the equations of motion. As a first seed for the Newton-Raphson algorithm, we choose the reference metric with $y_{0}=1/10$ and then march our code for increasing or decreasing values of $y_{0}$. Our selection of reference metric has restricted our temperature range to $T\,L>3/2^{4/3}\approx 1.19$; this is (A.10) with $y_{0}=0$. Our solutions have temperature above this critical value. To get the Helmoltz free energy $\mathcal{F}^{(0)}$ we integrate the first law of thermodynamics, $\mathrm{d}\mathcal{F}^{(0)}=-\mathcal{S}^{(0)}_{H}\,\mathrm{d}\mathcal{T}^{(0)}_{H}$, and the energy is then $\mathcal{E}^{(0)}=\mathcal{F}^{(0)}+\mathcal{T}^{(0)}_{H}\mathcal{S}^{(0)}_{H}$. The tension $\mathcal{T}_{Z}^{(0)}$ along the circle $S^{1}_{L}$ can be obtained from the Smarr relation (12), $8\,\mathcal{E}^{(0)}=9\,\mathcal{T}_{H}^{(0)}\,\mathcal{S}_{H}^{(0)}+\mathcal{T}_{Z}^{(0)}$. The values of these thermodynamics quantities are in excellent agreement with the results obtained using the standadard Arnowitt-Deser-Misner formalism [49, 71] or the covariant Noether charge formalism (a.k.a. covariant phase space method) [46, 47].
@@ A.2 Non-uniform strings
where the $\ldots$ has two types of contributions. First, we have higher order polynomial contributions of the form $(1-x)^{n}$ with $n\geq 10$. Second, there are nonperturbative terms in 1x1-x of the form $e^{-\frac{2\pi\,n}{1-x}}\cos(2\pi ny)$ (for $i=1,2,4,5$) and $e^{-\frac{2\pi\,n}{1-x}}\sin(2\pi ny)$ (for $i=3$). The coefficients $\delta_{i}$ are independent of $y$ and will later parametrise the energy and tension of the non-uniform string.